OpenStudy (anonymous):

How to solve cot 11 pi/4

OpenStudy (anonymous):

\[\cot(\frac{ 11\pi }{ 4 })\] \[=\cot(\frac{ 11\times180 }{ 4 })\] \[=\cot495\] \[=\cot(495-360)\] \[=\cot(135)\]

OpenStudy (anonymous):

how would i turn it into tan though, i know how to do the rest.

OpenStudy (anonymous):

\[\cot\theta =\frac{ 1 }{ \tan\theta } \]

OpenStudy (anonymous):

So will I just have to sub the degree and find its value or will it be the same degree of tan and cot

OpenStudy (anonymous):

I don't get what you're trying to say because you're telling me two options that are practically the same.

OpenStudy (anonymous):

\[\cot135=\frac{ 1 }{ \tan135 }\]

OpenStudy (anonymous):

Use your calculator to find the value.